OpenStudy (anonymous):

Let F(x)=2x+2,g(x)=x^2, and h(x)=3/x-2 State the domain of the operation. F(x)+g(x)

OpenStudy (anonymous):

F(x)+g(x) = 2x+2 + x^2 = x^2 +2x+2 domain is all real numbers